# Curriculum effects and compositionality emerge with in-context learning in neural networks

### v3

@@ Abstract
Human learning embodies a striking duality: sometimes, we appear capable of following logical, compositional rules and benefit from structured curricula (e.g., in formal education), while other times, we rely on an incremental approach or trial-and-error, learning better from curricula that are unstructured or randomly interleaved. Influential psychological theories explain this seemingly disparate behavioral evidence by positing two qualitatively different learning systems—one for rapid, rule-based inferences and another for slow, incremental adaptation. It remains unclear how to reconcile such theories with neural networks, which learn via incremental weight updates and are thus a natural model for the latter type of learning, but are not obviously compatible with the former. However, recent evidence suggests that both metalearning neural networks and large language models are capable of “in-context learning” (ICL)—the ability to flexibly grasp the structure of a new task from a few examples given at inference time. Here, we show that networks capable of ICL can reproduce human-like learning and compositional behavior on rule-governed tasks, while at the same time replicating human behavioral phenomena in tasks lacking rule-like structure via their usual in-weight learning (IWL). Our work shows how emergent ICL can equip neural networks with fundamentally different learning properties than those traditionally attributed to them, and that these can coexist with the properties of their native IWL, thus offering a novel perspective on dual-process theories and human cognitive flexibility.
1 Introduction
@@
Humans are capable of two qualitatively distinct kinds of learning [1, 2, 3, 4, 5, 6, 7, 8, 9]. The first involves slow, incremental adaptation to the environment through trial and error [2, 9, 6]. The second is much more advanced and involves rapid inference of rules or structure from information available in the environment or held in working memory [10, 11, 12, 13, 14]. For example, although it can famously take 10,000 hours to master the violin, when given a mandolin for the first time an expert musician may rapidly infer the rules about how each string is tuned.
@@
Many findings from cognitive psychology and neuroscience support the idea that humans exhibit different learning and generalization behaviors in different domains [1, 15, 5, 16]. In tasks that are readily described by simple rules (e.g., tasks where a feature like color determines correct responses), humans learn efficiently from only a few examples, appearing to make rapid inferences about the latent structure governing the task [12, 17, 18]. They also appear capable of generalizing this structure compositionally, flexibly recombining familiar elements into novel combinations according to the inferred rules [19, 20, 21, 22, 23, 24, 25, 26, 27]. In such settings, people exhibit a blocking advantage, learning better when information is organized into blocks of related examples that make this underlying structure more salient [28, 19, 15, 5]. In contrast, when a task is not governed by simple rules, learning may require integrating across multiple task dimensions, and proceeds much more incrementally [1, 29, 5]. In these contexts, compositional generalization is not possible, and people show an interleaving advantage, learning better when trials are randomly shuffled over time in both laboratory [5, 30] and real-world contexts [31, 32].
@@
Dual-process accounts [1, 4, 33, 5], explain these contrasting effects by positing two separate learning systems: a rule-based or symbolic system that is compositional and operates by testing explicit hypotheses, and a procedural or sub-symbolic system that learns more incrementally and can capture arbitrary associations, even in the absence of simple rules. Neural networks offer a natural framework for understanding the latter: standard networks operate by incrementally updating their weights [34, 35] and exhibit catastrophic forgetting when learning experiences are blocked, but not when they are interleaved [36, 37, 38]. However, it is less clear how they could explain the blocking advantage exhibited by humans on some tasks [15, 5]. Moreover, neural networks have traditionally been criticized for failing to account for human compositionality [20, 24, 39, 40, 41], as they do not explicitly represent rules or symbols [42, 12, 43]. Some biologically informed neural network models account for various aspects of dual-process theories [e.g., 44], such as models of prefrontal cortex (PFC) that emphasize the importance of dynamic activation-based representations for inferring rules and flexibly adapting to the current context [10, 45, 11, 46, 14, 47]. However, these models have not confronted how the emergence of such rule-based processing might relate to curriculum effects and compositionality.
@@
Figure 1: Properties of in-context learning (ICL) and in-weight learning (IWL). ICL (blue) is the ability of a neural network to flexibly learn a new task from just a few study examples given in context, and to apply this knowledge to a novel test query (also given in context). ICL is carried out through the activation dynamics of the network (blue arrow), and can happen without weight updates. ICL can be compositional, and is shown here predicting the location of a blue alligator (x=3, y=2) by composing elements of the known locations of a blue elephant (x=3) and a red alligator (y = 2). ICL also exhibits a blocking advantage, learning better when related examples are blocked over time. IWL (red) is the usual form of learning in neural networks, wherein prediction errors are backpropagated to update weights. IWL is non-compositional, depicted here as failing to generalize to the blue alligator due to its reliance on a simple lookup table that ignores the compositional structure of the task. IWL exhibits an interleaving advantage, learning better when examples are randomly shuffled or interleaved due to the well-known problem of catastrophic forgetting.
@@
We hypothesized that both compositionality and human curriculum effects might be reproduced by neural networks with greater cognitive flexibility. Recent advances in artificial neural networks have demonstrated surprising success on rule-governed tasks involving reasoning [48, 49, 50], analogy [51, 52], and compositionality [53, 54, 55]. Many of these capabilities are connected to the emergence of in-context learning (ICL), or the ability to learn new tasks from demonstrations or instructions given in context [48, 56, 57, 58]. For example, if demonstrations of a novel task are provided as contextual inputs (strawberry → red, banana → yellow), pretrained networks such as large language models (LLMs) can often readily perform the task on new inputs (plum → ??).
@@
Importantly, ICL does not require updates to network weights. This stands in contrast with in-weight learning (IWL)—the usual form of learning in neural networks—which proceeds by backpropagating errors to update weights [34]. Instead, ICL takes place within the model’s activation dynamics, similar to the dynamics supporting working memory in neural network models of PFC [13, 14, 10, 59]. Thus, the emergence of a capacity for ICL results in a tradeoff [60, 56, 61, 62]: when ICL succeeds, fewer errors are accumulated, resulting in fewer updates to the weights. This tradeoff resembles one seen in humans, wherein working memory facilitates rapid acquisition of stimulus-response rules but suppresses prediction errors, leading to degraded reinforcement learning and less robust retention [9, 63].
@@
Advanced ICL abilities have been shown to emerge in LLMs [48, 49], but can also be imparted more directly via metalearning, where a network is specifically trained to learn how to learn new tasks provided in context [64, 65, 66, 59, 47]. Metalearning networks that perform ICL through their activation dynamics have been shown to reproduce phenomena associated with the PFC [14, 10, 47], and human-like compositional generalizations [53], suggesting that emergent ICL algorithms can be more rule-like or compositional than the standard IWL used to train networks in the first place [26, 67].
@@
In this work, we demonstrate how a single neural network capable of both ICL and IWL can simultaneously replicate the behavioral effects associated with each of the two systems posited in traditional theories [1, 4, 5], producing compositional generalization and the blocking advantage in rule-governed tasks, while exhibiting an interleaving advantage in tasks lacking such structure. Our theoretical framework can be summarized by three key principles (see Figure 1):
@@
    Standard IWL fails on compositional generalization problems, and shows an interleaving advantage due to catastrophic forgetting when trials are blocked.
@@
    ICL can be endowed with inductive biases that produce compositional generalization and a blocking advantage.
@@
    When ICL is possible, its properties dominate because few errors are made and IWL is suppressed. But when ICL is difficult, the learning properties of IWL dominate because more errors are made and backpropagated.
@@
We test this theoretical framework by experimenting with metalearning neural networks on tasks based on those used in previous human studies [5, 19]. First, we show in a category-learning setting [5] that a single neural network capable of ICL and IWL produces both of the curriculum effects observed in humans—a blocking advantage in the presence of rule-like structure, and an interleaving advantage in the absence of such structure. Then, we show that when applied to a compositional task, the neural network produces the compositional generalization behaviors and the blocking advantage observed in humans on the same task [19]. Finally, we test existing pretrained LLMs on this compositional task and show that their emergent ICL algorithms exhibit both compositionality and a blocking advantage. Taken together, our findings show how two qualitatively distinct learning processes can coexist in a single neural network model, thus offering a unique perspective on how a dual-process architecture might emerge in a neural network.
2 Results
2.1 Curriculum effects in category-learning
@@
We first consider whether the principles above can account for the curriculum effects observed in human category learning, before turning to compositionality in the next section. As reviewed above, in category learning, humans exhibit a blocking advantage when categories are governed by succinct rules, but an interleaving advantage when no such rules are readily available [5].
@@
Figure 2: Category-learning experiments. The task is derived from a human study [5]. Networks were presented with multi-feature items along with their category labels, and tested on unseen items. See Appendix for details. (a) Curriculum conditions. Trials were either blocked by category label or randomly interleaved. (b) In the rule-like condition, category membership was determined by a simple rule that only depended on one of the two features (e.g., ‘A’ if length $\leq$ 4, ‘B’ otherwise). (c) In the rotated condition, category membership was determined by both feature dimensions. The original axes were rotated by 45 degrees and a category boundary was chosen in the new coordinate system. (d-f) Category-learning with in-weight learning (IWL) only. Randomly initialized networks were trained from scratch on the task. (d) The few-shot evaluation tested networks’ ability to learn the task from the 32 examples presented in context, before any weight updates were made. Unsurprisingly, randomly initialized networks without prior metalearning experience were incapable of utilizing the examples given in context to learn the task, regardless of condition. Values correspond to the average test accuracies shown in (f), but before step 0 (i.e., before any finetuning took place). (e) The same evaluation was conducted after “finetuning”, testing networks’ ability to learn the task through IWL. Without prior metalearning experience, the network was able to learn in-weights, performing well on both the rule-like and the rotated tasks after training. However, performance was much worse in the blocked condition due to catastrophic forgetting (see f). Here, values correspond to the train accuracy (i.e., accuracy on the 32 train items) in (f) at the final timestep. (f) Accuracy and loss results over the course of IWL training in each of the four conditions. Accuracy is split by category to better visualize the effects of catastrophic forgetting in the blocked condition (top row). (g-i) Category learning with both in-weight and in-context learning (ICL). Networks first metalearned on a distribution of related tasks (not shown), and were subsequently finetuned on specific category-learning tasks from each condition. (g) After metalearning, the models exhibited strong ICL on the task, as shown by the high few-shot accuracy. ICL demonstrated a blocking advantage, and also showed improved performance in the rule-like compared to the rotated condition. (h) After finetuning, the network exhibited an interleaving advantage in the rotated condition, due to catastrophic forgetting when trials were blocked (see i). (i) Accuracy and loss results over the course of finetuning in each of the four conditions. When trials were blocked in the rule-like condition, ICL achieved near-perfect accuracy immediately, resulting in little loss and thus little IWL. When trials were interleaved, few-shot performance was worse (see g), but performance quickly recovered due to compensation by IWL. In the rotated condition, ICL failed, resulting in larger losses and increased IWL. This IWL resulted in catastrophic forgetting, as can be seen in the rapid decline in performance on ‘A’ items while training on ‘B’, and vice versa. No such catastrophic forgetting occurred when trials were interleaved (although test performance was not perfect).
@@
We designed a category-learning task directly based on this previous work [5], but suitable for use with metalearning neural networks (see Figure 2a-c). Stimuli varied along two feature dimensions (akin to line length and line orientation) with 8 possible values, yielding 64 possible items. Each item was assigned to one of two categories, indicated by an arbitrary category label (e.g., ‘A’ or ‘B’). In the Rule-like condition, one of the two feature dimensions determined category membership (e.g., lines with shorter lengths are in category ‘A’ and lines with longer lengths are in category ‘B’), while in the Rotated condition, category membership was determined by both features. This simple rotation has been shown to challenge the search for a simple, verbalizable rule, and is thought to recruit the more incremental procedural learning system in humans [5, 1]. Networks were presented with 16 items from each category (32 total), and tested on the remaining held-out items. The 32 items used during learning were either Blocked, where items from one category were presented first, followed by the items from the other, or Interleaved, where items were randomly shuffled. Both rotation conditions were tested with both curriculum conditions, yielding a 2x2 design.
2.1.1 IWL produces an interleaving advantage
@@
2.1.2 ICL can produce a blocking advantage
Next, we endowed a network with ICL abilities by (meta)training it on a distribution of categorization tasks (see methods for details). Metalearning can induce ICL in deep neural networks [53, 57], and relatedly, has been shown to give rise to abstract generalizable representations in models of PFC [14, 47, 10]. These ICL abilities allowed the network to solve unseen tasks given in context through its activation dynamics, even when weights were frozen and no IWL was allowed to occur.
To ensure that the emergent ICL algorithm would have the desired properties (see principle 2), we (meta)trained it on a distribution of categorization tasks with 1) rule-like structure and 2) blocked curricula. We then evaluated the trained network in the few-shot setting, where the weights were frozen and the network had to learn new tasks from a few examples given in context (see methods for details; see Figure 2g). As predicted, when the model was endowed with an ICL algorithm familiar with rule-like category-learning problems, it could easily generalize to new rule-like problems, but struggled to solve tasks in context in the rotated condition (main effect of rotation: $p<10^{-3}$). Moreover, the emergent ICL algorithm exhibited a blocking advantage on unseen rule-like categorization problems (main effect of curriculum: $p<0.05$). This blocking advantage emerged due to the metalearning distribution (see Appendix), but see Discussion for alternative explanations based on architectural constraints in human brains. In sum, these few-shot results suggest that it is possible to endow a network with an ICL algorithm that is sensitive to rule-like structure and to the learning curriculum: the network’s forward activation dynamics were capable of making inferences over the items provided in context, but was better at doing so when related items were blocked over time.
2.1.3 Concurrent ICL and IWL reproduce both curriculum effects
@@
As we described above, when categories are governed by rule-like structure, ICL succeeds on the task and exhibits a blocking advantage in few-shot inference. But in the rotated task, where categories are not governed by rule-like structure, ICL struggles (Figure 2g). The resulting errors, when backpropagated, drive an increase in IWL, producing an interleaving advantage due to catastrophic forgetting (Figure 2i; interaction between curriculum and rotation: $p<10^{-3}$).
@@
Thus, consistent with our predictions, we have shown that a single model capable of ICL and IWL can recapitulate the curriculum effects observed in human category-learning [5]. When the network is capable of making inferences over familiar rules, it can solve new tasks from a few samples given in context. However, when the environment does not afford such inferences or the network cannot make them, IWL can still compensate, allowing good performance. This IWL suffers from catastrophic forgetting, resulting in an interleaving advantage on the rotated task.
2.2 Curriculum effects in a compositional task
@@
As noted above, one of the most impressive recent developments in neural networks has been the demonstration that ICL can give rise to compositionality [53, 55, 67], traditionally considered to be a major theoretical challenge to neural networks [20, 40]. Recent results have shown that while standard IWL in neural networks struggles to reproduce human-like compositional generalization behaviors [39, 68, 69], emergent ICL abilities in neural networks can appear to compose inferred rules in order to generalize to new inputs [49, 53, 54, 55]. Thus, a key goal of our framework is to leverage the distinction between ICL and IWL to provide a unified account of both the compositional generalization behaviors and the curriculum effects observed in humans. In particular, ICL should account for both the blocking advantage and for compositional generalization in tasks governed by rule-like structure, while IWL accounts for the interleaving advantage observed when such compositional generalization is challenging or impossible.

### v4

# The dynamic interplay between in-context and in-weight learning in humans and neural networks

@@ Abstract
Human learning embodies a striking duality: sometimes, we appear capable of following logical, compositional rules and benefit from structured curricula (e.g., in formal education), while other times, we rely on an incremental approach or trial-and-error, learning better from curricula that are randomly interleaved. Influential psychological theories explain this seemingly disparate behavioral evidence by positing two qualitatively different learning systems—one for rapid, rule-based inferences and another for slow, incremental adaptation. It remains unclear how to reconcile such theories with neural networks, which learn via incremental weight updates and are thus a natural model for the latter type of learning, but are not obviously compatible with the former. However, recent evidence suggests that metalearning neural networks and large language models are capable of “in-context learning” (ICL)—the ability to flexibly grasp the structure of a new task from a few examples. Here, we show that the dynamic interplay between ICL and default in-weight learning (IWL) naturally captures a broad range of learning phenomena observed in humans, reproducing curriculum effects on category-learning and compositional tasks, and recapitulating a tradeoff between flexibility and retention. Our work shows how emergent ICL can equip neural networks with fundamentally different learning properties that can coexist with their native IWL, thus offering a novel perspective on dual-process theories and human cognitive flexibility.
Introduction
@@
Humans are capable of two qualitatively distinct kinds of learning [1, 2, 3, 4, 5, 6, 7, 8, 9]. The first involves slow, incremental adaptation to the environment through trial and error [2, 9, 6]. The second is much more advanced and involves rapid inference of rules or structure from information available in the environment or held in working memory [WM; 10, 11, 12, 13, 14]. For example, although it can famously take 10,000 hours to master the violin, when given a mandolin for the first time an expert musician may rapidly infer the rules about how each string is tuned.
@@
Many findings support the idea that humans exhibit different learning and generalization behaviors in different domains [1, 15, 5, 16, 17, 11, 9]. On the one hand, in tasks that are readily described by simple rules (e.g., where only one stimulus feature is relevant), humans learn efficiently from only a few examples (“few-shot learning”), appearing to make rapid inferences about the latent structure governing the task [12, 18, 19]. When this latent structure is compositional, humans can generalize by flexibly recombining familiar elements according to inferred rules [20, 21, 22, 23, 24, 25, 26, 27, 28]. In such settings, people exhibit a blocking advantage, learning better when information is organized into blocks of related examples that make this underlying structure more salient [29, 20, 15, 5]. On the other hand, when a task is not governed by simple rules, learning may require integrating across multiple task dimensions. This kind of learning proceeds more incrementally [1, 30, 5], but is also associated with greater retention after a delay [9, 17]. In these contexts, compositional generalization is not possible, and, as shown in both laboratory [5, 31] and real-world contexts [32, 33], people exhibit an interleaving advantage, learning better when trials are randomly shuffled over time.
@@
Dual-process accounts [1, 4, 34, 5], explain these contrasting effects by positing two separate learning systems: a rule-based or symbolic system that is compositional and operates by testing explicit hypotheses, and a procedural or sub-symbolic system that learns more incrementally and can capture arbitrary associations, even in the absence of simple rules. Neural networks, which learn via incremental weight updates [35, 36], offer a natural framework for understanding the latter. However, it is less clear how they could explain the features associated with the rule-based system, namely, flexible (few-shot) learning, compositionality, and the blocking advantage. Neural networks are notoriously data hungry compared to human learners [12], and have been criticized for failing to account for human compositionality [21, 25, 37, 38, 39], as they do not explicitly represent rules or symbols [40, 12, 41]. Furthermore, the blocked curricula that confer an advantage to humans on some tasks [15, 5] generally result in “catastrophic forgetting” in neural networks, because new learning can overwrite information stored in the same weights during previous blocks [42, 43, 44].
Some biologically informed models account for various aspects of dual-process theories [e.g., 45], such as models of prefrontal cortex (PFC) that emphasize the importance of dynamic activation-based representations for inferring rules and flexibly adapting to the current context [10, 46, 47, 48, 14, 49]. However, these models have not confronted how the emergence of such rule-based processing might relate to curriculum effects and compositionality.
@@
Figure 1: Properties of in-context learning (ICL) and in-weight learning (IWL). ICL (blue) is the ability of a neural network to flexibly learn a new task from just a few study examples given in context, and to apply this knowledge to a novel test query (also given in context). ICL is carried out through the activation dynamics of the network (blue arrow), and can happen without weight updates. ICL can be flexible and compositional, and is shown here predicting the location of a blue alligator (x=3, y=2) by composing elements of the known locations of a blue elephant (x=3) and a red alligator (y = 2). ICL can exhibit a blocking advantage, learning better when related examples are blocked over time. IWL (red) is the usual form of learning in neural networks, wherein prediction errors are backpropagated to update weights. IWL can result in better retention but is non-compositional, depicted here as failing to generalize to the blue alligator due to its reliance on a simple lookup table that ignores the compositional structure of the task. IWL exhibits an interleaving advantage, learning better when examples are randomly shuffled or interleaved due to the well-known problem of catastrophic forgetting.
@@
We hypothesized that neural networks with greater cognitive flexibility can simultaneously reproduce human-like behaviors related to few-shot generalization, compositionality and curriculum effects. Recent advances in artificial neural networks have demonstrated surprising success on rule-governed tasks involving reasoning [50, 51, 52], analogy [53, 54], and compositionality [55, 56, 57]. Many of these capabilities are connected to the emergence of in-context learning (ICL), or the ability to learn new tasks from demonstrations or instructions given in context [50, 58, 59, 60]. For example, if demonstrations of a novel task are provided as contextual inputs (strawberry → red, banana → yellow), trained networks such as large language models (LLMs) can often readily perform the task on new inputs (plum → ??).
@@
Importantly, ICL does not require updates to network weights. [^1] One may prefer to think of ICL not as a kind of “learning,” but rather as a kind of inference. This terminological issue can also arise in the case of human few-shot learning. We have chosen to adopt the “learning” terminology as it has become standard in machine learning research. This stands in contrast with in-weight learning (IWL)—the usual form of learning in neural networks—which proceeds by backpropagating errors to update weights [35]. Instead, ICL takes place within the model’s activation dynamics, which similarly support cognitive flexibility and working memory (WM) in neural network models of PFC [13, 14, 10, 61]. Thus, the emergence of ICL results in a tradeoff [62, 58, 63]: when ICL succeeds, fewer errors are accumulated, resulting in fewer weight updates. We reasoned that such a tradeoff could account for a variety of phenomena in human learning. For example, when information can be learned rapidly within WM, neural prediction errors are suppressed [9, 64, 11]. This neural signature predicts enhanced generalization of rule-like structure [11] but degraded reinforcement learning and less robust retention [17, 65].
@@
Advanced ICL abilities have been shown to emerge in LLMs [50, 51], but can also be imparted more directly via metalearning, where a network is specifically trained (via the usual form of IWL) to learn how to learn new tasks provided in context [66, 67, 68, 69, 61, 49]. Metalearning networks that perform ICL through their activation dynamics have been shown to reproduce phenomena associated with the PFC [14, 10, 49], and human-like compositional generalizations [55], suggesting that emergent ICL abilities can result in more rule-like or compositional behaviors than the standard IWL used to train networks in the first place [27, 70].
@@
In this work, we demonstrate how a single neural network capable of both ICL and IWL can simultaneously replicate the behavioral effects associated with each of the two systems posited in traditional dual-process theories [1, 4, 5], producing compositional generalizations and the blocking advantage in rule-governed tasks, while exhibiting an interleaving advantage in tasks lacking such structure. Moreover, we show how the very same mechanisms can give rise to the tradeoff between flexibility and retention observed in human reinforcement learning [17, 11, 65].
Our theoretical framework can be summarized by three key principles (see Figure 1):
@@
    Properties of IWL: IWL fails on compositional generalization problems, shows an interleaving advantage due to catastrophic forgetting when trials are blocked, and results in better retention.
@@
    Properties of ICL: ICL can be endowed with inductive biases that produce compositional generalization and a blocking advantage, but results in worse retention.
@@
    Dynamic interplay: when ICL is possible, its properties dominate because few errors are made, suppressing IWL. But when ICL is difficult, the properties of IWL dominate because errors result in larger weight updates.
@@
We test this theoretical framework by experimenting with metalearning neural networks on tasks used in previous human studies [5, 20], showing how the dynamic interplay between ICL and IWL offers a unified account of learning phenomena observed in humans across a wide range of studies from cognitive psychology and neuroscience [^2] All code used for simulations is available at https://github.com/jlrussin/icl-iwl-interplay.. First, we show in a category-learning setting [5] that a single neural network capable of ICL and IWL produces both of the curriculum effects observed in humans—a blocking advantage in the presence of rule-like structure, and an interleaving advantage in the absence of such structure. Then, we show that when applied to a compositional task, the neural network produces the compositional generalization behaviors and the blocking advantage observed in humans on the same task [20]. We then test existing LLMs on this compositional task and show that emergent ICL in these models exhibits both compositionality and a blocking advantage as well. Finally, we demonstrate how the dynamic interplay between ICL and IWL naturally gives rise to a tradeoff between flexibility and retention observed in recent human studies [9, 17, 65].
The primary goal of this work is not simply to show that neural networks can perform well on these cognitive tasks, but to demonstrate how the specific principles governing the dynamic interplay between ICL and IWL naturally reproduce these particular learning phenomena. Our findings show how these two qualitatively distinct learning processes can interact within a single neural network model, and offer a novel perspective on how dual-process theories of cognition might be reconciled with a neural network perspective.
Results
Curriculum effects in category-learning
@@ Results
We first consider whether the principles above can account for the curriculum effects observed in human category learning, before turning to compositionality in the next section. As reviewed above, humans exhibit an interaction between the task (rule-like vs. rotated), and curriculum (blocked vs. interleaved) in category learning ($\eta_{p}=0.04$), showing a blocking advantage ($d=0.47$) when categories are governed by succinct rules, but an interleaving advantage ($d=0.33$) when no such rules are readily available [5].
@@
Figure 2: Category-learning experiments. The task is derived from a human study [5]. Networks were presented with multi-feature items along with their category labels, and tested on unseen items. See Appendix for details. (a) Curriculum conditions. Trials were either blocked by category label or randomly interleaved. (b) In the rule-like condition, category membership was determined by a simple rule that only depended on one of the two features (e.g., ‘A’ if length $\leq$ 4, ‘B’ otherwise). (c) In the rotated (information-integration) condition, category membership was jointly determined by both feature dimensions. The original axes were rotated by 45 degrees and a category boundary was chosen in the new coordinate system. (d-f) Category-learning with in-weight learning (IWL) only. Randomly initialized networks were trained from scratch on the task. (d) The few-shot evaluation tested networks’ ability to learn the task from the 32 examples presented in context, before any weight updates were made. Unsurprisingly, randomly initialized networks without prior metalearning experience were incapable of utilizing the examples given in context to learn the task, regardless of condition. Values correspond to the average test accuracies shown in (f), but before step 0 (i.e., before any training). Note that model choices were not forced to be one of the two category labels, so chance performance here corresponds to $1/d_{v}$, where $d_{v}$ is the vocabulary size. (e) The same evaluation was conducted after training, testing networks’ ability to learn through IWL. Without prior metalearning experience, the network was able to learn in-weights, performing well on both the rule-like and the rotated tasks after training. However, performance was much worse in the blocked condition due to catastrophic forgetting (see f). Here, values correspond to the train accuracy (i.e., accuracy on the 32 train items) in (f) at the final timestep. (f) Accuracy and loss results over the course of IWL training in each of the four conditions. Accuracy is split by category to better visualize the effects of catastrophic forgetting in the blocked condition (top row). (g-i) Category learning with both in-weight and in-context learning (ICL). Networks first metalearned on a distribution of related tasks (not shown), and were subsequently trained on specific category-learning tasks from each condition. (g) After metalearning, the models exhibited strong ICL on the task, as shown by the high few-shot accuracy. ICL demonstrated a blocking advantage, and also showed improved performance in the rule-like compared to the rotated condition. (h) After training had occurred on a specific task, the network exhibited an interleaving advantage in the rotated condition, due to catastrophic forgetting when trials were blocked (see i). (i) Accuracy and loss results over the course of task-specific training in each of the four conditions. When trials were blocked in the rule-like condition, ICL achieved near-perfect accuracy immediately, resulting in little loss and thus little IWL. When trials were interleaved, few-shot performance was worse (see g), but performance quickly recovered due to compensation by IWL. In the rotated condition, ICL failed, resulting in larger losses and increased IWL. This IWL resulted in catastrophic forgetting, as can be seen in the rapid decline in performance on ‘A’ items while training on ‘B’, and vice versa. No such catastrophic forgetting occurred when trials were interleaved (although test performance was not perfect).
@@
We designed a category-learning task directly based on this previous work [5], but suitable for use with metalearning neural networks (see Figure 2a-c). Stimuli varied along two feature dimensions (akin to line length and line orientation) with 8 possible values, yielding 64 possible items. Each item was assigned to one of two categories, indicated by an arbitrary category label (e.g., ‘A’ or ‘B’). In the Rule-like (or “Rule-based”) condition, one of the two feature dimensions determined category membership (e.g., lines with shorter lengths are in category ‘A’ and lines with longer lengths are in category ‘B’), while in the Rotated (or “information-integration”) condition, category membership was determined by both features. This simple rotation has been shown to challenge the search for a simple, verbalizable rule, and is thought to recruit the more incremental procedural learning system in humans [5, 1]. Networks were presented with 16 items from each category (32 total), and tested on the remaining held-out items. The 32 items used during learning were either Blocked, where items from one category were presented first, followed by the items from the other, or Interleaved, where items were randomly shuffled. Both rotation conditions were tested with both curriculum conditions, yielding a 2x2 design.
IWL produces an interleaving advantage
@@
ICL can produce a blocking advantage
Next, we endowed a network with ICL abilities by having it metalearn on a distribution of categorization tasks (see methods for details). Metalearning can induce ICL in deep neural networks [55, 59], and relatedly, has been shown to give rise to abstract generalizable representations in models of PFC [14, 49, 10]. These ICL abilities allowed the network to solve unseen tasks given in context through its activation dynamics, even when weights were frozen and no IWL was allowed to occur.
To ensure that ICL would have the desired properties (see principle 2), we had the network metalearn on a distribution of categorization tasks with 1) rule-like structure and 2) blocked curricula. We then evaluated the network in the few-shot setting, where the weights were frozen and the network had to learn new tasks from a few examples given in context (see methods for details; see Figure 2g). As predicted, when the model had developed ICL abilities by metalearning on rule-like category-learning problems, it could generalize to new rule-like problems, but struggled to solve tasks in context in the rotated condition (main effect of rotation: $p<10^{-3}$). Moreover, ICL exhibited a blocking advantage on unseen rule-like categorization problems (main effect of curriculum: $p<0.05$). This blocking advantage also emerged due to the metalearning distribution (see Appendix), but see Discussion for alternative explanations based on architectural constraints in human brains. In sum, these few-shot results suggest that it is possible to endow a network with ICL abilities that are sensitive to rule-like structure and learning curriculum: the network was capable of making inferences over the items provided in context, but was better at doing so when related items were organized into blocks.
Concurrent ICL and IWL reproduce both curriculum effects
@@
As we described above, when categories are governed by rule-like structure, ICL succeeds on the task and exhibits a blocking advantage in few-shot inference. But in the rotated task, where categories are not governed by rule-like structure, ICL struggles (Figure 2g). The resulting errors drive an increase in IWL, producing an interleaving advantage due to catastrophic forgetting (Figure 2i; interaction between curriculum and rotation: $p<10^{-3}$).
@@
Taken together, the above experiments show that a single model capable of ICL and IWL can recapitulate the curriculum effects observed in human category-learning [5]. When the network is capable of making inferences over familiar rules, it can solve new tasks from a few samples given in context. However, when the environment does not afford such inferences or the network cannot make them, IWL can still compensate, allowing good performance. This IWL suffers from catastrophic forgetting, resulting in an interleaving advantage on the rotated task.
Curriculum effects in a compositional task
@@
As noted above, one of the most impressive recent developments in research on neural networks has been the demonstration that ICL can give rise to compositionality [55, 57, 70], traditionally considered to be a major theoretical challenge to neural networks [21, 38]. Recent studies have shown that while standard IWL struggles to reproduce human-like compositional generalization behaviors [37, 71, 72], ICL can appear to compose inferred rules in order to generalize to new inputs [51, 55, 56, 57]. Thus, a key goal of our framework is to leverage the distinction between ICL and IWL to provide a unified account of both the compositional generalization behaviors and the curriculum effects observed in humans. In particular, ICL should account for both the blocking advantage and the compositional generalization behaviors observed in tasks governed by rule-like structure, while IWL accounts for the interleaving advantage observed when such compositional generalization is challenging or impossible.